# LabUtopia: High-Fidelity Simulation and Hierarchical Benchmark for Scientific Embodied Agents

### v1

@@ Abstract
Scientific embodied agents play a crucial role in modern laboratories by automating complex experimental workflows. Compared to typical household environments, laboratory settings impose significantly higher demands on perception of physical-chemical transformations and long-horizon planning, making them an ideal testbed for advancing embodied intelligence. However, its development has been long hampered by the lack of suitable simulator and benchmarks. In this paper, we address this gap by introducing LabUtopia, a comprehensive simulation and benchmarking suite designed to facilitate the development of generalizable, reasoning-capable embodied agents in laboratory settings. Specifically, it integrates i) LabSim, a high-fidelity simulator supporting multi-physics and chemically meaningful interactions; ii) LabScene, a scalable procedural generator for diverse scientific scenes; and iii) LabBench, a hierarchical benchmark spanning five levels of complexity from atomic actions to long-horizon mobile manipulation. LabUtopia supports 30 distinct tasks and includes more than 200 scene and instrument assets, enabling large-scale training and principled evaluation in high-complexity environments. We demonstrate that LabUtopia offers a powerful platform for advancing the integration of perception, planning, and control in scientific-purpose agents and provides a rigorous testbed for exploring the practical capabilities and generalization limits of embodied intelligence in future research.
@@ 1 Introduction
A promising approach to address the above issues is the sim-to-real framework zhao2020sim , where agents are first trained within realistic simulations before being deployed in real-world laboratory settings. This paradigm enables cost-effective and safe training while maintaining the potential for generalization to physical environments hsu2023sim . However, current established simulators predominantly focus on household environments kolve2017ai2 ; puig2018virtualhome ; yang2024holodeck and fail to adequately address the specific challenges of scientific experimentation. As discussed in Table 1, they exhibit three fundamental limitations: (1) the inability to model chemical dynamics, such as product formation or color changes, which are essential for accurate perception and reasoning in lab tasks; (2) limited diversity and semantic richness in their asset libraries, which restricts the faithful representation of heterogeneous laboratory environments in real world; and (3) the lack of comprehensive evaluation protocols, particularly those that span from fine-grained atomic actions to complex, long-horizon experimental procedures.
@@ 3.1 LabSim: High-Fidelity Simulation Environment
Physical Realism. LabSim supports physically accurate interactions among rigid, deformable, and fluid entities PhysX5 . Each asset in the environment is annotated with empirically grounded physical properties such as mass, friction coefficients, and restitution. For rigid-body simulation, we enable precise contact and collision modeling. For soft materials, we incorporate deformable body physics to capture compressible and elastic behavior. Notably, for fluid simulation, we employ a GPU-accelerated Position-Based Dynamics (PBD) framework macklin2013position , supporting rich fluid-agent interactions required for scientific manipulation.
Chemical Process Modeling. To simulate chemical processes within laboratory tasks, we introduce a chemical engine that integrates a curated knowledge base with a reasoning model. We begin by constructing a structured database of 200 common chemical substances, sourced from the authoritative PubChem repository kim2016pubchem . Each encodes key its attributes, such as color, molar mass, and pH value, allowing it to be represented as a substance asset within the simulation. Given a set of reactants, we leverage a large language model (GPT-4o mini hurst2024gpt ) to reason about potential chemical processes and infer corresponding transformations, including color changes, product formation, etc. These inferred changes are then rendered in the simulation by dynamically updating the physical state and visual properties of the involved substances. This engine equips LabUtopia with the capability to model complex chemical interactions with both interpretability and flexibility.
3.2 LabScene: Procedural Laboratory Environment Generation at Scale
@@
Scene Assets. Due to the scarcity of open-source lab assets in the community, we collected over one thousand candidate scenes from designer websites. These raw assets underwent a multi-stage pre-processing pipeline involving content filtering, format normalization, and structural standardization. To ensure realism, we consulted domain experts in chemistry and physics to assess scene fidelity and provide refinement suggestions. Based on their feedback, we selected a high-quality subset of expert-verified scenes to serve as the foundational environments for LabUtopia.
Object Assets. While some collected scenes include basic instruments, many lack detailed internal structures required for executing lab tasks. To address this, we assembled a comprehensive set of object assets spanning essential equipment (e.g., drying ovens) and diverse glassware types (e.g., beakers). To ensure compatibility with robotic manipulation, we refine and modularize these assets into standard, interactable objects. And some were augmented with articulated joints and hinge mechanisms to support physically plausible interactions within the simulation. In total, the final dataset includes about 60 categories of laboratory equipment assets and around 80 types of transparent glassware and plasticware assets, covering a diverse range of materials, sizes, and functional forms.
Environment Generation Pipeline. To incorporate our collected instruments into the laboratory scenes, we develop an environment generation pipeline that balances layout diversity with physical plausibility. Specifically, all objects are placed sequentially based on importance and size rankings. For each object, candidate positions and orientations are sampled from a discretized grid that satisfies various constraints, including boundary, collision, and instrument-specific constraints deitke2022️ . The layout score is computed considering factors such as edge proximity, inter-object distance, and orientation alignment. The configuration with the highest score is selected. If the random sampling fails to produce a valid layout within a time limit, the system falls back to a depth-first search strategy sun2024forest2seq , systematically exploring placements while enforcing physical and spatial constraints. This hybrid approach ensures functional, reasonable scene layouts suitable for embodied agent training.
@@ 3.3 Task-Rich Trajectory Collection
Manipulation Trajectory Auto-collection. We divide the motion planners into two levels: atomic action controllers and task-level action controllers. Atomic actions are standard laboratory operations, such as pouring and stirring, that align with experimental protocols and are executed using finite state machines. Task-level controllers organize atomic action controllers to collect data for specific tasks. For atomic actions, we control multiple target keypoints for the robotic arm. At each keypoint, an RMPflow controller cheng2020rmp plans motion toward dynamically determined positions based on the real-time state of manipulated objects. We incorporate spherical linear interpolation (Slerp) gong2023arnold for continuously manipulating articulated objects, e.g., opening the dry container task, for more robust results. Furthermore, task-level controllers organize atomic actions to streamline the entire experimental procedure, enabling our motion planner to generate demonstration data efficiently.
@@ 4.2 Evaluation Protocol
Embodiment. We employ a 7-DoF Franka Emika Panda FrankaRobotics2025 manipulator with a parallel gripper for manipulation tasks. The agent controls the seven joints and gripper directly. Isaac Sim’s built-in motion planner maps end-effector commands to joint-space actions for execution. For navigation tasks, we utilize the Fetch mobile manipulator FetchRobotics2025 and Ridgeback with Franka ClearpathRobotics2025 , both supporting arm manipulation and base locomotion. These robots are controlled using three degrees of freedom—x and y velocities and rotation—enabling integrated navigation and manipulation wu2020towards .
Evaluation Execution. A task instance is considered successful if the success condition—where the current state remains within the tolerance threshold of the goal state—is continuously satisfied for 2 seconds after the agent completes the final stage of the task, which is the same as previous work gong2023arnold ; he2024learning ; hu2024videopredictionpolicygeneralist . For example, in the “Open Door” task, success requires the cabinet door to remain at the specified open position for 2 seconds after the robot releases the handle, ensuring no shortcuts during motion. The success rate is used as the evaluation metric in LabSim, with strict evaluation ensuring both the robot and object maintain the successful state for the required duration, only after the motion planner has executed its final action.
@@ 5.1 Experimental Setup
Models. To benchmark the performance of existing imitation learning algorithms in LabSim, we select two representative models: ACT fu2024mobile and Diffusion Policy chi2023diffusion .
• ACT is a transformer-based model designed for sequential decision making in robotic manipulation. ACT processes RGB images, robot proprioception, and tokenized instructions through a multi-layer transformer encoder. At each decision step, ACT autoregressively predicts the next low-level action, conditioned on past observations and actions.
• Diffusion Policy is a generative model that formulates robot control as a conditional diffusion process. The model takes recent observations as input and learns to generate control trajectories by progressively denoising an initial random trajectory sample, conditioned on the observation context. This approach not only enables efficient learning from expert demonstrations, but also provides robust stochasticity during action generation. In our work, we utilize the CNN-based version of Diffusion Policy as the encoder.
Table 2: Performance comparison across task levels. For Level-3, results are reported under ID/OOD settings. Values represent success rates (%).
@@
Level-3 Pick 81.7 / 71.7 53.3 / 41.7
Press 98.3 / 96.7 81.7 / 31.6
Open Door 73.3 / 65.0 63.3 / 58.3
Pour Liquid 75.0 / 65.0 46.6 / 31.6
Heater Beaker 86.7 / 80.0 21.6 / 8.3
Transport Beaker 77.5 / 73.3 67.5 / 15.0
Visual Input. We use two cameras for each manipulation task, positioned to ensure both the robot and the objects are visible. Camera placement may vary slightly by task, but the general principle remains consistent. Each camera outputs 256×256 RGB images by default. Since accurate depth data for transparent objects is difficult to obtain in the real world, depth maps are not included by default, but users can render images at any resolution and optionally enable depth maps. Other Omniverse sensors (e.g., tactile sensors) are currently disabled as they are unnecessary for current tasks, but full support is available if needed.
Training Details. All experiments are conducted on a single Nvidia 4090 GPU. For each task, the collected dataset consists of 150 episodes. Evaluation is performed in real-time within a simulated environment based on the task settings, with 60 episodes used for testing. Following the original settings in ACT and Diffusion Policy, we set the horizon to 8 and 60, respectively, and the number of training epochs is uniformly set to 200.
@@
Level-4 Task Model Success Rate (%)
@@
5.2 Experimental Results
We evaluated the performance of the ACT and DP models across Level-1 to Level-3 tasks in Table 2. For Level-1 tasks, both models demonstrated robust fundamental manipulation skills, achieving high success rates due to the simplicity of these single-step operations. In Level-2 tasks, success rates began to decline, particularly for DP. We observed that DP frequently stalled. For instance, in the Heater Beaker task, the DP algorithm halted during the final action until time expired, while in the Operate Drawer task, most DP failures stemmed from its inability to execute any action successfully. In the Stir with GlassRod task, both models exhibited errors in grasping or stirring position due to the small size of the glass rod.
Task Level Model Success Rate (%)
ID OOD
Pick ACT 31.2 1.7
DP 11.2 0.0
Pour Liquid ACT 25.6 0.0
DP 8.0 0.0
In Level-3 tasks, the performance gap widened, with ACT maintaining superior stability, while DP struggled with generalization to novel materials and precise actions, such as pressing a button in Heater Beaker. In contrast, ACT showed greater robustness when handling out-of-domain visual features. We tested generalization to out-of-domain shapes in Table 4 by jointly training datasets with objects of varying sizes and evaluating on out-of-domain objects. The results revealed that joint training with objects of significantly different sizes led to a substantial drop in in-domain success rates, with near-zero success rates for out-of-domain shapes and sizes. This suggests that both models largely lack the ability to manipulate out-of-domain objects. Level-4 tasks, as shown in the provided table, assessed long-horizon tasks. ACT significantly outperformed DP, achieving higher success rates in single-stage primitive (SP) actions and sub-steps (A1–A7), though both models experienced sharp declines in later sub-steps (A5–A7) due to cumulative errors in complex sequences. In sub-steps, although each policy is more stable when handling its own task, it is necessary to design specific action decomposition and switching mechanisms, which increases the overall system complexity. Moreover, the transitions between actions may become discontinuous or suffer from distribution shift issues.

### v2

@@ Abstract
Scientific embodied agents play a crucial role in modern laboratories by automating complex experimental workflows. Compared to typical household environments, laboratory settings impose significantly higher demands on perception of physical-chemical transformations and long-horizon planning, making them an ideal testbed for advancing embodied intelligence. However, its development has been long hampered by the lack of suitable simulator and benchmarks. In this paper, we address this gap by introducing LabUtopia, a comprehensive simulation and benchmarking suite designed to facilitate the development of generalizable, reasoning-capable embodied agents in laboratory settings. Specifically, it integrates i) LabSim, a high-fidelity simulator supporting multi-physics and chemically meaningful interactions; ii) LabScene, a scalable procedural generator for diverse scientific scenes; and iii) LabBench, a hierarchical benchmark spanning five levels of complexity from atomic actions to long-horizon mobile manipulation. LabUtopia supports 30 distinct tasks and includes more than 200 scene and instrument assets, enabling large-scale training and principled evaluation in high-complexity environments. We demonstrate that LabUtopia offers a powerful platform for advancing the integration of perception, planning, and control in scientific-purpose agents and provides a rigorous testbed for exploring the practical capabilities and generalization limits of embodied intelligence in future research. Project web page: the github URL.
@@ 1 Introduction
A promising approach to address the above issues is the sim-to-real framework zhao2020sim, where agents are first trained within realistic simulations before being deployed in real-world laboratory settings. This paradigm enables cost-effective and safe training while maintaining the potential for generalization to physical environments hsu2023sim; li2025chemistry3d; team2025robobrain. However, current established simulators predominantly focus on household environments kolve2017ai2; puig2018virtualhome; zhou2025code; yang2024holodeck; kang2025viki; zhou2025roborefer and fail to adequately address the specific challenges of scientific experimentation. As discussed in Table 1, they exhibit three fundamental limitations: (1) the inability to model chemical dynamics, such as product formation or color changes, which are essential for accurate perception and reasoning in lab tasks; (2) limited diversity and semantic richness in their asset libraries, which restricts the faithful representation of heterogeneous laboratory environments in real world; and (3) the lack of comprehensive evaluation protocols, particularly those that span from fine-grained atomic actions to complex, long-horizon experimental procedures.
@@ 3.1 LabSim: High-Fidelity Simulation Environment
Physical Realism. LabSim supports physically accurate interactions among rigid, deformable, and fluid entities PhysX5. Each asset in the environment is annotated with empirically grounded physical properties, we enable precise contact and collision modeling. For soft materials, we incorporate deformable body physics to capture compressible and elastic behavior. Notably, for fluid simulation, we employ a GPU-accelerated Position-Based Dynamics (PBD) framework macklin2013position, supporting rich fluid-agent interactions required for scientific manipulation.
Chemical Process Modeling. To simulate chemical processes within laboratory tasks, we introduce a chemical engine that integrates a curated knowledge base with a reasoning model. We begin by constructing a structured database of 200 common chemical substances, sourced from the authoritative PubChem repository kim2016pubchem. Each encodes its key attributes, such as color, molar mass, and pH value, allowing it to be represented as a substance asset within the simulation. Given a set of reactants, we leverage a large language model (GPT-4o hurst2024gpt) to reason about potential chemical processes and infer corresponding transformations, including color changes, product formation, etc. These inferred changes are then rendered in the simulation by dynamically updating the physical state and visual properties of the involved substances. This engine equips LabUtopia with the capability to model complex chemical interactions with both interpretability and flexibility.
3.2 LabScene: A Large-Scale Dataset of Scientific Laboratory Scenes and Instruments
@@
Scene Assets. Due to the scarcity of open-source lab assets in the community, we collected a large number of candidate scenes from designer websites. These raw assets underwent a multi-stage pre-processing pipeline, including content filtering, format normalization, and structural standardization. To ensure realism, we consulted experts in chemistry and physics to assess the fidelity of the scenes and provide refinement suggestions. Based on their feedback, we selected approximately 100 high-quality, expert-verified scenes to serve as the foundational environments for LabUtopia.
Object Assets. While some of the collected laboratory scenes already include basic instruments and furnishings, many lack the fine-grained internal structures and detailed geometry necessary for accurately simulating and executing real-world laboratory tasks. To overcome this limitation, we curated and constructed a comprehensive library of high-fidelity object assets that represent a wide range of laboratory tools and apparatus. These assets encompass essential equipment such as drying ovens, centrifuges, pipettes, and balances, as well as diverse glassware and plasticware types including beakers, flasks, test tubes, and Petri dishes. To ensure full compatibility with robotic manipulation and physics-based simulation, all assets were refined, standardized, and modularized into physically consistent, interactable forms. The final collection consists of approximately 60 categories of laboratory equipment and over 80 types of transparent glassware and plasticware, covering a rich diversity of materials, scales, and functional configurations.
Environment Generation Pipeline. To incorporate our collected instruments into the laboratory scenes, we develop an environment generation pipeline that balances layout diversity with physical plausibility. Specifically, all objects are placed sequentially based on importance and size rankings. For each object, candidate positions and orientations are sampled from a discretized grid that satisfies various constraints, including boundary, collision, and instrument-specific constraints deitke2022️; yang2025llm. The layout score is computed considering factors such as edge proximity, inter-object distance, and orientation alignment. The configuration with the highest score is selected. If the random sampling fails to produce a valid layout within a time limit, the system falls back to a depth-first search strategy sun2024forest2seq, systematically exploring placements while enforcing physical and spatial constraints. This hybrid approach ensures functional, reasonable scene layouts suitable for embodied agent training.
@@ 3.3 Task-Rich Trajectory Collection
Manipulation Trajectory Auto-collection. We divide the motion planners into two levels: atomic action controllers and task-level action controllers. Atomic actions are standard laboratory operations, such as pouring and stirring, that align with experimental protocols and are executed using finite state machines. Task-level controllers organize atomic action controllers to collect data for specific tasks. For atomic actions, we control multiple target keypoints for the robotic arm. At each keypoint, an RMPflow controller cheng2020rmp plans motion toward dynamically determined positions based on the real-time state of manipulated objects. We incorporate spherical linear interpolation (Slerp) gong2023arnold for continuously manipulating articulated objects, e.g., opening the dry box task, for more robust results. Furthermore, task-level controllers organize atomic actions to streamline the entire experimental procedure, enabling our motion planner to generate demonstration data for imitation learning efficiently. During demonstration data collection, all objects are randomly initialized within a predefined spatial range to ensure diversity and generalization.
@@
Figure 3: (a) Workspace for manipulation tasks. (b) Illustration of the navigation environment. (c) Bird’s-eye view of the navigation map. (d) Occupancy grid map used for navigation. (f) Occupancy grid map with the planned path highlighted. ‘
@@ 4.2 Evaluation Protocol
Embodiment. We employ a 7-DoF Franka Emika Panda FrankaRobotics2025 manipulator with a parallel gripper for manipulation tasks. For navigation tasks, we utilize the Fetch mobile manipulator FetchRobotics2025 and a mobile manipulation robot composed of the Clearpath Robotics Ridgeback base integrated with a Franka Emika Panda arm ClearpathRobotics2025, both supporting arm manipulation and base locomotion. These robots are controlled using three degrees of freedom—x and y velocities and rotation—enabling integrated navigation and manipulation wu2020towards.
Evaluation Execution. Object positions are randomly placed within a 15 cm × 15 cm region, adjusted according to their sizes and workspace constraints to ensure compatibility with task requirements. A task instance is considered successful if the current state remains within the tolerance threshold of the goal state and continuously satisfied for 2 seconds after completing the final task stage, consistent with prior works gong2023arnold; he2024learning; hu2024videopredictionpolicygeneralist; liu2024robomamba. For example, in the “Open Door” task, success requires the cabinet door to remain at the specified open position for 2 seconds after the robot releases the handle, ensuring no shortcuts during motion. The success rate is used as the evaluation metric in LabBench, with strict evaluation ensuring both the robot and object maintain the successful state for the required duration, only after the motion planner has executed its final action. More details about the task descriptions, visualizations, and success determination criteria can be found in Appendix C.
@@ 5.1 Experimental Setup
Table 2: Performance comparison across task levels. Values represent success rates (%).
@@
Models. To benchmark the performance of existing imitation learning algorithms in LabSim, we select two representative models: ACT fu2024mobile, Diffusion Policy chi2023diffusion, and $\pi_{0}$ black2024pi_0
• ACT is a transformer-based model designed for action chunk prediction in robotic manipulation. ACT processes RGB images, robot proprioception through a multi-layer transformer encoder. At each decision step, ACT autoregressively predicts the next low-level action, conditioned on past observations and actions.
• Diffusion Policy is a generative model that formulates robot control as a conditional diffusion process. The model takes recent observations as input and learns to generate control trajectories by progressively denoising an initial random trajectory sample, conditioned on the observation context. In our work, we utilize the CNN-based version of Diffusion Policy.
$\pi_{0}$ is a vision-language-action (VLA) model based on PaliGemma, which enables modeling of complex continuous actions by incorporating an additional action expert output module and utilizing flow matching techniques.
Model Input And Output.
All manipulation tasks utilize three cameras (a wrist camera on the robotic arm, one facing the robotic arm, and one directly overhead), with each camera outputting 256×256 RGB images by default. DP and ACT models use images from the overhead and front-facing cameras, while $pi_{0}$ model uses images from the wrist camera and the camera facing the robotic arm. Users can also render images at any resolution or output depth maps. Other Omniverse sensors (e.g., tactile sensors) are currently disabled as they are unnecessary for the current tasks, but full support is available if needed. The input for all models also includes the absolute values of the robot’s joint positions. The model’s output is the predicted joint positions of the robotic arm for the next time step.
Training Details Our training parameters are primarily adapted from the open-source ACT, Diffusion Policy and OpenPI algorithms. ACT and Diffusion Policy were implemented using PyTorch, with training and evaluation conducted on a single NVIDIA RTX 4090 GPU. $\pi_{0}$ was implemented using JAX, with training on 8 A800 80G GPUs. ACT and Diffusion Policy were trained for 200 epochs using the AdamW optimizer and Exponential Moving Average (EMA). $\pi_{0}$ was trained for 30,000 steps. Model inputs include two camera views, each with a resolution of $256\times 256$ pixels, as well as robot joint position inputs. The output consists of predicted joint parameters. A fixed learning rate of $1\times 10^{-4}$ was used throughout. The batch sizes were set to 128 for ACT and 64 for Diffusion Policy. Additionally, ACT and $\pi_{0}$ were trained using a single frame of historical observation, whereas Diffusion Policy was trained using three consecutive frames as input. The action chunk lengths for ACT, DP, and $\pi_{0}$ are 60, 60, and 8, respectively.
Table 3: Performance comparison across Level-3 task. Results are reported under ID/OOD settings. Values represent success rates (%).
Task Level Task Name $\bm{\pi_{0}}$ ACT DP
Level-3 Pick 83.3/85.8 81.7 / 71.7 53.3 / 41.7
Press 92.5/89.1 98.3 / 96.7 81.7 / 31.6
Open Door 51.6/53.3 73.3 / 65.0 63.3 / 58.3
Pour Liquid 40.0/38.3 75.0 / 65.0 46.6 / 31.6
Heater Beaker 89.1/86.7 86.7 / 80.0 21.6 / 8.3
Transport Beaker 86.7/88.3 77.5 / 73.3 67.5 / 15.0
5.2 Experimental Results
We evaluated the performance of the ACT and DP models across Level-1 to Level-3 tasks in Table 2 and Table 3, and additionally tested the $\pi_{0}$ model on Level-3 tasks. For Level-1 tasks, both models demonstrated robust fundamental manipulation abilities, achieving high success rates due to the simplicity and single-step nature of these tasks. In Level-2 tasks, the success rates began to decline, particularly for the DP model. We observed that DP frequently stalled during execution. For example, in the Heater Beaker task, the DP algorithm paused during the final action until the time limit expired; while in the Operate Drawer task, most DP failures resulted from its inability to successfully execute any action. In the Stir with GlassRod task, both models exhibited noticeable errors in grasping or stirring positions due to the small size of the glass rod.
Task Level Model Success Rate (%)
ID OOD
Pick ACT 31.2 1.7
DP 11.2 0.0
Pour Liquid ACT 25.6 0.0
DP 8.0 0.0
In Level-3 tasks, the performance gap widened, with ACT maintaining superior stability, while DP struggled with generalization to novel materials and precise actions, such as pressing a button in Heater Beaker. In contrast, ACT showed greater robustness when handling out-of-domain visual features. Additionally, we evaluated the VLA-based model $\pi_{0}$ on selected Level-3 generalization tasks. While pretrained VLA models, after fine-tuning, exhibited minimal performance degradation on out-of-distribution visual inputs or novel materials, they did not consistently outperform models trained from scratch in these tasks. We tested generalization to out-of-domain shapes in Table 4 by jointly training datasets with objects of varying sizes and evaluating on out-of-domain objects. The results revealed that joint training with objects of significantly different sizes led to a substantial drop in in-domain success rates, with near-zero success rates for out-of-domain shapes and sizes. This suggests that both models largely lack the ability to manipulate out-of-domain objects.
In Level-4 tasks, as shown in Table 5, we assessed the performance in long-horizon tasks. ACT significantly outperformed DP, achieving higher success rates in single-stage primitive (SP) actions and sub-steps (A1–A7), though both models experienced sharp declines in later sub-steps (A5–A7) due to cumulative errors in complex sequences. In sub-steps, although each policy is more stable when handling its own task, it is necessary to design specific action decomposition and switching mechanisms, which increases the overall system complexity. Moreover, the transitions between actions may become discontinuous or suffer from distribution shift issues.
@@ 5.2 Experimental Results
Level-4 Task
Model
Success Rate (%)